### v1

Abstract.
The axiomatization of harmonic measure theory is established. As the applications, Dirichlet boundary problem in graph theory and Dahlberg’s theory are generalized under the generalized harmonic measure theory. The theory provides an interpretation from the measure theory view point of the boundary value problem to the elliptic equation.
Key words and phrases:
Harmonic measure; Harnonic function; Strong $(p,p)$-inequality.
1991 Mathematics Subject Classification:
28A75, 31A15, 31C05
1. Introduction
1.1. The classical harmonic measure
It is well known that harmonic measure can be defined in several analytic ways, one of which has relationship with generalized Dirichlet problem, see [1, 3]. The generalized Dirichlet problem is to find a harmonic function $h$ on $D$ corresponding to the boundary function $f$, where $D$ is any nonempty open subset of $\mathbb{R}^{n}$ and $f$ is any extended real-valued function on $\partial D$.
In order to give the basic concept of harmonic measure, we need some related conceptions. An extended real-valued function $u$ is called hyperharmonic (hypoharmonic) on $D$ if it is superharmonic (subharmonic) or identically +$\infty$+ ∞ ($\infty$- ∞) on each component of $D$. Denote
[…]
$\mathop{\lim\inf}\limits_{y\rightarrow x}u(y)\geq f(x),\forall x\in\partial D\},$
[…]
$\mathop{\lim\sup}\limits_{y\rightarrow x}u(y)\leq f(x),\forall x\in\partial D\},$
and $\overline{H}_{f}:=\inf\{u:u\in\mathcal{U}_{f}\}$, $\underline{H}_{f}:=\sup\{u:u\in\mathcal{L}_{f}\}$. If $\overline{H}_{f}=\underline{H}_{f}$ and both harmonic on $D$, then $f$ is called resolutive and $H_{f}:=\overline{H}_{f}=\underline{H}_{f}$ is called the Dirichlet solution for $f$. The following theorem confirms that the continuous function is resolutive.
Theorem 1.1 (Wiener).
If $f$ is a continuous real-valued function on the boundary $\partial D$ of the bounded open set $D$, then $f$ is resolutive.
Suppose $D$ is a bounded open set. Denote $L_{x}(f):=H_{f}(x)$ for any $f\in C(\partial D)$ and any $x\in D$. Then for each $x\in D$, $L_{x}(f)$ is a positive functional on $C(\partial D)$ and moreover, by the Riesz representation theorem, there is a unique unit measure $\omega_{x}$ on the Borel subsets of $\partial D$ such that $H_{f}(x)=L_{x}(f)=\int f\,\mathrm{d}\omega_{x}$. The measure $\omega_{x}$ on $C(\partial D)$ is called the harmonic measure relative to $D$ and $x$.
Our aim in the current paper is to give a self-contained harmonic measure theory in the framework of measure theory in Hausdorff spaces. In this theory, Maximum Principle holds in open set depending on the regular closeness of the harmonic system. Harnack Principle holds depending on the translation and scaling invariant of the harmonic system and the boundedness of the relative kernel. Some applications are presented in this paper, including a generalization of graph theory and Dahlberg’s theory.
The paper is organized as follows. In Section 1, we give the classical definition of the harmonic measure in an analytic way. In Section 2, the axiomatization of harmonic measure is presented, including the Maximal principle, Harnack’s inequality and some related topics. In Section 3, a generalization of the Dirichlet boundary problem in the graph theory is considered. In Section 4, a generalization of Dahlberg’s theory is considered.
2. The axiomatization of harmonic measure
In this section, we give the basic axiomatization of the harmonic measure.
2.1. Topological space theory
We list some basic concepts and conclusions of topological space theory.
Suppose $X$ is a nonempty set. The pair $(X;\tau)$, or briefly denoted by $X$, is called a topological space with the open sets family $\tau\subseteq 2^{X}$ provided
  1. (1)
    $X,\emptyset\in\tau$;
  2. (2)
    $\mathop{\cup}\limits_{\lambda\in\Lambda}O_{\lambda}\in\tau$ for any $\{O_{\lambda}\}_{\lambda\in\Lambda}\subseteq\tau$;
  3. (3)
    $O_{\alpha}\cap O_{\beta}\in\tau$ for any $O_{\alpha},O_{\beta}\in\tau$.
Elements of $\tau$ are called the open sets of $X$. Any open set containing $x\in X$ is called a neighborhood of $x$ and is denoted by $O_{x}$. Let $\{x_{n}\}_{n\in\mathbb{N}}\subseteq X$. By $x_{n}\rightarrow x_{0}$ we mean for any open set $U\ni x_{0}$, there exists an $N\in\mathbb{N}$, for any $n>N$, such that $x_{n}\in U$. Please note that in some topological spaces, for $x_{0}\not=y_{0}$, both $x_{n}\rightarrow x_{0}$ and $x_{n}\rightarrow y_{0}$ may happen at the same time. We say $x\in X$ is a boundary point of the subset $U\subseteq X$, if any neighborhood of $x$ both contains both a point of $U$ and a point of $U^{c}:=X\backslash U$. All boundary points of $U$ is denoted by $\partial U$. A subset $U\subseteq X$ is always considered as a topological space $(U;\tau_{U})$ with the open sets family $\tau_{U}:=\{U\cap O:O\in\tau\}$. We say that $A$ is a compact set in $X$ if any open cover of $A$ has a finite subcover.
Let $X$ and $Y$ be two topological spaces. A map $f:X\rightarrow Y$ is called continuous if for each open set $U\subseteq Y$, the set $f^{-1}(U)$ is open.
We say a topological space is a Hausdorff space if for any $x,y\in X\,(x\not=y)$, there exist separately two neighborhood of $x,y$, denoted by $O_{x},O_{y}\in\tau$, such that $O_{x}\cap O_{y}=\emptyset$. In a Hausdorff space, if $x_{n}\rightarrow x_{0}$, then $x_{0}$ must be unique.
2.2. Measure theory
We list some basic concepts and conclusions of measure theory.
Suppose $X$ is a nonempty set. A mapping $\omega:2^{X}\rightarrow[0,\infty]$ is called a measure on $X$ provided
  1. (1)
    $\omega(\emptyset)=0$;
  2. (2)
    if $A\subseteq\mathop{\cup}\limits_{k=1}^{\infty}A_{k}$, then $\omega(A)\leq\sum\limits_{k=1}^{\infty}\omega(A_{k})$.
For any $A\subseteq X$ the mapping $\delta:2^{X}\rightarrow\{0,1\}$ defined by
$\delta_{x}(A)=\left\{\begin{aligned} 1,\text{ if }x\in A;\\ 0,\text{ if }x\not\in A,\end{aligned}\right.$
is a measure on $X$. We always call this measure Dirac measure.
A set $A\subseteq X$ is $\omega$-measurable if for each set $B\subseteq X$ such that
$\omega(B)=\omega(B\cap A)+\omega(B\backslash A).$
If $\{A_{k}\}_{k=1}^{\infty}$ is a sequence of disjoint $\omega$-measurable sets, then
$\omega\left(\mathop{\cup}\limits_{k=1}^{\infty}A_{k}\right)=\sum_{k=1}^{\infty}\omega(A_{k}).$
A measure $\omega$ on $X$ is called Borel measure if each Borel set in $X$ is $\omega$-measurable.
Assume $\mu$ and $\nu$ are measures on $X$. The measure $\nu$ is called absolutely continuous with respect to $\mu$, written $\nu\ll\mu$, provided for any $A\subseteq X$, $\mu(A)=0$ implies $\nu(A)=0$; The measure $\nu$ and $\mu$ are called mutually singular, written $\nu\perp\mu$, if there exists a $B\subseteq X$ such that $\mu(X\backslash B)=\nu(B)=0$.
Let $X$ be a set and $Y$ be a topological space. A map $f:X\rightarrow Y$ is called $\omega$-measurable if for each open set $U\subseteq Y$, the set $f^{-1}(U)$ is $\omega$-measurable. A function $f:X\rightarrow[-\infty,+\infty]$ is $\omega$-measurable if and only if $f^{-1}([-\infty,a))$ is $\omega$-measurable for any $a\in\mathbb{R}$.
A function $g:X\rightarrow[-\infty,+\infty]$ is called a simple function if the image of $g$ is finite. Denote $g^{\pm}:=\max\{\pm g,0\}$. For simple, $\omega$-measurable function $g$, if either $\sum\limits_{0\leq y\leq\infty}y\omega((g^{+})^{-1}(y))<\infty$ or $\sum\limits_{0\leq y\leq\infty}y\omega((g^{-})^{-1}(y))<\infty$, we define its integral
$\int g\,\mathrm{d}\omega:$ $=\int g^{+}\,\mathrm{d}\omega-\int g^{-}\,\mathrm{d}\omega$
$:=\sum_{0\leq y\leq\infty}y\omega((g^{+})^{-1}(y))-\sum_{0\leq y\leq\infty}y\omega((g^{-})^{-1}(y)),$
[…]
$\inf\left\{\int g\,\mathrm{d}\omega:g\text{ is simple, }\omega\text{-measurable},g\geq f,\int g\,\mathrm{d}\omega\text{ exists}\right\}$
=:absent:=:= : $\int^{*}f\,\mathrm{d}\omega=\int_{*}f\,\mathrm{d}\omega$
::=:= $\sup\left\{\int g\,\mathrm{d}\omega:g\text{ is simple, }\omega\text{-measurable},g\leq f,\int g\,\mathrm{d}\omega\text{ exists}\right\}.$
It is well known that the integral of a function $f:X\rightarrow[-\infty,+\infty]$ exists if and only if $f$ is $\omega$-measurable. A function $f:X\rightarrow[-\infty,+\infty]$ is called $\omega$-integrable if $f$ is $\omega$-measurable and $\int|f|\,\mathrm{d}\omega<\infty$. For $\delta_{x}$-measurable function $f$, $\int f\,\mathrm{d}\delta_{x}=f(x)$.
For $\omega$-measurable function we always denote $\omega(f):=\int f\,\mathrm{d}\omega$, and for measurable set $A\subseteq X$, briefly denote $\omega(A):=\omega(\chi_{A})=\int\chi_{A}\,\mathrm{d}\omega$.
For $1\leq p<\infty$, $L^{p}(X;\omega)$ (resp. $L_{\text{loc}}^{p}(X;\omega)$) denotes the set of all $\omega$-measurable functions on $X$ such that $|f|^{p}$ is integrable (resp. locally integrable). $L^{\infty}(X;\omega)$ (resp. $L_{\text{loc}}^{\infty}(X;\omega)$) denotes the set of all $\omega$-measurable functions on $X$ such that $|f|$ is essentially bounded (resp. locally and essentially bounded).
Let $\omega,\omega_{k}\,(k\in\mathbb{N})$ be measures on topological space $X$. We say the measures $\{\omega_{k}\}_{k\in\mathbb{N}}$ weakly converge to the measure $\omega$, written $\omega_{k}\rightharpoonup\omega$ if for all $f\in C_{c}(X)$,
$\lim_{k\rightarrow\infty}\omega_{k}(f)=\omega(f),$
where $C_{c}(X)$ consists of continuous functions on $X$ with compact support.
2.3. The axiomatization of harmonic measure
We give the definition of harmonic measure defined on the Hausdorff space.
In this article, we always suppose $X$ is a Hausdorff space.
Definition 2.1 (Pre-harmonic measure).
Suppose $D\subset X$ is open, with $\partial D\not=\emptyset$. A family of Borel measures $\{\omega^{D}_{x}:2^{\partial D}\rightarrow[0,1]:\,x\in D\}$ on $\partial D$, briefly denoted by $\{\omega^{D}_{x}\}$, is called a pre-harmonic measure relative to $D$ provided
  1. (1)
    For any $x\in D$, $\omega^{D}_{x}(\partial D)=1$;
  2. (2)
    For any $x,y\in D$, $\omega^{D}_{x}\ll\omega^{D}_{y}$.
From (2) we get the following proposition.
Proposition 2.1.
Suppose $\{\omega^{D}_{x}\}$ is a pre-harmonic measure relative to $D$. If $W$ is a component of $D$, then the class of $\omega^{D}_{x}$-measure 00 Borel subsets of $\partial D$ is independent of $x\in W$.
For different $x$ and $y$ in $D$, the measurable sets and functions of $\omega_{x}^{D}$ and $\omega_{y}^{D}$ may be separately different. We want to find common measurable sets and functions of $\{\omega_{x}^{D}:x\in D\}$. Denote
[…]
Then $\mathcal{F}_{x}$ is a $\sigma$-algebra. Let $\mathcal{F}:=\mathop{\cap}\limits_{x\in D}\mathcal{F}_{x}$. Then $\mathcal{F}$ is a $\sigma$-Algebra containing Borel subsets of $\partial D$. Since $\omega^{D}_{x}$ can be uniquely extended to $\mathcal{F}_{x}$, the same is true of $\mathcal{F}$. The extension of $\omega^{D}_{x}$ to $\mathcal{F}$ is also denoted by $\omega^{D}_{x}$.
Definition 2.2 (Harmonic measure).
(1) Elements in $\mathcal{F}$ are called $\omega_{\cdot}^{D}$-measurable sets and extended real-valued functions measurable relative to $\mathcal{F}$ are called $\omega_{\cdot}^{D}$-measurable functions. The pre-harmonic measure $\{\omega_{x}^{D}\}$ restricted on $\mathcal{F}$, briefly denoted by $\omega_{\cdot}^{D}$, is called the harmonic measure relative to $D$.
(2) A $\omega_{\cdot}^{D}$-measurable function $f$ on $\partial D$ is called $\omega_{\cdot}^{D}$-integrable if it is integrable relative to $\omega_{x}^{D}$ for any $x\in D$.
(3) A harmonic measure $\omega_{\cdot}^{D}$ is called continuous in $D$ if $h_{f}(x):=\omega_{x}^{D}(f)$ is continuous in $D$ for any $f\in L(\partial D;\omega_{\cdot}^{D})$.
(4) A point $y\in\partial D$ is called a regular point relative to $\omega_{\cdot}^{D}$ if for any $f\in L(\partial D;\omega_{\cdot}^{D})$ continuous at $y\in\partial D$ such that
(2.1) $\lim_{D\ni x_{n}\rightarrow y}\omega_{x_{n}}^{D}(f)=f(y).$
$\omega_{\cdot}^{D}$ or $D$ is called regular if every point of $\partial D$ is regular relative to $\omega_{\cdot}^{D}$.
For any $f\in L(\partial D;\omega_{\cdot}^{D})$ and any $y\in\partial D$, we always denote $\omega_{y}^{D}(f):=f(y)$.
We list several simple properties in the following proposition.
Proposition 2.2.
(1) $C(\partial D)\subseteq L(\partial D;\omega_{\cdot}^{D})$;
(2) Suppose $W$ is a component of $D$. Then the class of $\omega^{D}_{x}$-integrable functions is independent of $x\in W$.
(3) Suppose $y\in\partial D$ is a regular point. Then $\omega_{x_{n}}^{D}\rightharpoonup\delta_{y}$ as $D\ni x_{n}\rightarrow y$, that is, for any $f\in C(\partial D)$, such that
(2.2) $\lim_{D\ni x_{n}\rightarrow y}\omega_{x_{n}}^{D}(f)=f(y).$
(4) Suppose $y\in\partial D$ is a regular point. Then for any open set $V\ni y$ of $\partial D$, such that $\lim\limits_{D\ni x_{n}\rightarrow y}\omega_{x_{n}}^{D}(V)=1$.
Remark 2.1.
We only know the harmonic measure induced by the Laplace operator in $\mathbb{R}^{n}$ satisfies the equivalence of (2.1) and (2.2), see Corollary 8.21 in [3]. It is interesting to derive the condition to insure that (2.1) and (2.2) are equivalent in general topological spaces.
Proposition 2.3.
Suppose $\omega_{\cdot}^{D}$ is a harmonic measure. Then, for any nonempty open set $A\subseteq\partial D$ with a regular point in $A$, $\omega_{\cdot}^{D}(A)\not=0$.
Proof.
If there exists a nonempty open set $A$ with a regular point $y\in A$ such that $\omega_{\cdot}^{D}(A)=0$, then $\omega_{x}(\chi_{A})=\omega_{x}(A)=0$ for any $x\in D$. Since $\omega_{x}(\chi_{A})=\omega_{x}(A)=0$ is continuous with respect to $x\in D$, we get
$\lim_{D\ni x_{n}\rightarrow y}\omega_{x_{n}}(A)=0.$
At the same time, by Definition 2.2 (4), since $y\in A$ is regular and $\chi_{A}$ is continuous at $y\in A$, we get
$\lim_{D\ni x_{n}\rightarrow y}\omega_{x_{n}}(A)=\chi_{A}(y)=1,$
which contradicts to the above equation. ∎
Theorem 2.4 (Strong Maximum Principle for $\chi_{A}$ on regular open set).
Suppose $D$ is regular and $A\subseteq\partial D$ is open. Then
$\sup_{x\in D}\omega_{x}^{D}(\chi_{A})=\mathop{\max}\limits_{y\in\partial D}\chi_{A}(y),$
and if there exists an $x_{0}\in D$ such that
$\omega_{x_{0}}(\chi_{A})=\sup_{x\in D}\omega_{x}^{D}(\chi_{A}),$
then $\omega_{x}(\chi_{A})$ is constant for $x\in D$.
Proof.
When $A=\varnothing$, the conclusion holds clearly.
When $A\subseteq\partial D$ is a nonempty open set, by Proposition 2.3, we get
$0<\omega_{x}^{D}(\chi_{A})=\omega_{x}^{D}(A)\leq 1$
And for any $y\in A$, $y$ is regular and $\chi_{A}$ is continuous at $y\in A$, we get
$\lim_{D\ni x_{n}\rightarrow y}\omega_{x_{n}}(A)=\chi_{A}(y)=1.$
Then
$\sup_{x\in D}\omega_{x}^{D}(\chi_{A})=\mathop{\max}\limits_{y\in\partial D}\chi_{A}(y)=1.$
If there exists an $x_{0}\in D$ such that
$\omega_{x_{0}}(\chi_{A})=\sup_{x\in D}\omega_{x}^{D}(\chi_{A})=1,$
then $\omega_{x_{0}}(\partial D\backslash A)=0$. By Definition 2.1 (2) for any $x\in D$, $\omega_{x}(\partial D\backslash A)=0$. This implies that for any $x\in D$, $\omega_{x}(\chi_{A})=\omega_{x}(A)=1$. ∎
Theorem 2.5 (Weak Maximum Principle for regular harmonic measure).
Suppose $\omega_{\cdot}^{D}$ is regular. If $f\in C(\partial D)$ and there exists a $y_{0}\in\partial D$ such that $f(y_{0})=\mathop{\sup}\limits_{y\in\partial D}f(y)$, then
$\sup_{x\in D}\omega_{x}^{D}(f)=\mathop{\max}\limits_{y\in\partial D}f(y).$
Proof.
By (1) in Definition 2.1, for all $x\in D$, it is clear the flowing equation holds
$\omega_{x}^{D}(f)=\int_{\partial D}f\,\mathrm{d}\omega_{x}^{D}\leq\mathop{\max}\limits_{y\in\partial D}f(y)\int_{\partial D}\,\mathrm{d}\omega_{x}^{D}=\mathop{\max}\limits_{y\in\partial D}f(y).$
Then we get
$\sup_{x\in D}\omega_{x}^{D}(f)\leq\mathop{\max}\limits_{y\in\partial D}f(y).$
Since $f(y_{0})=\mathop{\max}\limits_{y\in\partial D}f(y)$, by Proposition 2.2 (3), $\lim\limits_{x_{n}\rightarrow y_{0}}\omega_{x_{n}}^{D}(f)=f(y_{0})$. Then
$\sup_{x\in D}\omega_{x}^{D}(f)=\mathop{\max}\limits_{y\in\partial D}f(y).$
Theorem 2.6 (Boundary Strong Maximum Principle).
Suppose $f\in L(\partial D;\omega_{\cdot}^{D})\cap L^{\infty}(\partial D;\omega_{\cdot}^{D})$. If there exists an $x_{0}\in D$ such that
$\omega_{x_{0}}(f)=\mathop{\mathrm{ess\,sup}}\limits_{y\in\partial D}f(y),$
then $\omega_{x}(f)$ is constant for $x\in D$.
Proof.
Denote $c_{0}=\mathop{\mathrm{ess\,sup}}\limits_{y\in\partial D}f(y)$, $A_{0}:=\{y\in\partial D:f(y)=c_{0}\}$, $A_{1}:=\{y\in\partial D:f(y)>c_{0}\}$ and $A_{2}:=\{y\in\partial D:f(y)<c_{0}\}$. It is clear that $\partial D=A_{0}\cup A_{1}\cup A_{2}$, $A_{i}(i=1,2,3)$ are disjoint and $\omega_{\cdot}^{D}(A_{1})=0$. We will prove $\omega_{\cdot}^{D}(A_{2})=0$. In fact by $\omega_{\cdot}^{D}(A_{1})=0$ we get
$c_{0}$ $=\omega_{x_{0}}(f)=\int_{\partial D}f\,\mathrm{d}\omega_{x_{0}}^{D}=\int_{\partial D\backslash A_{1}}f\,\mathrm{d}\omega_{x_{0}}^{D}$
$=\int_{\partial D\backslash A_{1}}[c_{0}+(f-c_{0})]\,\mathrm{d}\omega_{x_{0}}^{D}$
$=c_{0}+\int_{\partial D\backslash A_{1}}(f-c_{0})\,\mathrm{d}\omega_{x_{0}}^{D}.$
Then
$\int_{A_{2}}(f-c_{0})\,\mathrm{d}\omega_{x_{0}}^{D}=\int_{A_{2}\cup A_{0}}(f-c_{0})\,\mathrm{d}\omega_{x_{0}}^{D}=\int_{\partial D\backslash A_{1}}(f-c_{0})\,\mathrm{d}\omega_{x_{0}}^{D}=0.$
Since $f<c_{0}$ on $A_{2}$, we get $\omega_{\cdot}^{D}(A_{2})=0$. Then we can get
$\omega_{\cdot}^{D}(A_{0})=\omega_{\cdot}(\partial D)-\omega_{\cdot}^{D}(A_{1})-\omega_{\cdot}^{D}(A_{2})=1,$
which implies $f(y)=c_{0}$ for $\omega_{\cdot}^{D}$-a.e. $y\in\partial D$. Then for any $x\in D$
$\omega_{x}(f)=\int_{\partial D}f\,\mathrm{d}\omega_{x}=\int_{A_{0}}f\,\mathrm{d}\omega_{x}=c_{0}.$
By Theorem 2.5 and Theorem 2.6, we can get the following strong maximum principle.
Theorem 2.7 (Strong Maximum Principle for regular harmonic measure).
Suppose $\omega_{\cdot}^{D}$ is regular. If $f\in C(\partial D)$ and there exists a $y_{0}\in\partial D$ such that $f(y_{0})=\mathop{\sup}\limits_{y\in\partial D}f(y)$, then
$\sup_{x\in D}\omega_{x}^{D}(f)=\mathop{\max}\limits_{y\in\partial D}f(y),$
and if there exists an $x_{0}\in D$ such that
$\omega_{x_{0}}(f)=\sup_{x\in D}\omega_{x}^{D}(f),$
then $\omega_{x}(f)$ is constant for $x\in D$.
For different open sets in $X$, we need the following concept of harmonic system.
Definition 2.3 (Harmonic measure system).
(1) We call the family of continuous harmonic measures $\{\omega_{\cdot}^{D_{\lambda}}:\lambda\in\Lambda\}$ is a harmonic measure system on $X$, briefly denoted by $\{\omega_{\cdot}^{D_{\lambda}}\}$, if its regular elements are compatible, that is, for any $D\in\{D_{\lambda}:\lambda\in\Lambda\}$, any regular $\omega_{\cdot}^{\widetilde{D}}\in\{\omega_{\cdot}^{D_{\lambda}}\}$ with $\widetilde{D}\subseteq D$, any functions $f\in L(\partial D;\omega_{\cdot}^{D})$ and any $x\in\widetilde{D}$, such that $\omega^{D}_{x}(f)=\omega_{x}^{\widetilde{D}}(\omega_{\cdot}^{D}(f))$.
(2) In a harmonic measure system $\{\omega_{\cdot}^{D_{\lambda}}\}$, a point $y\in\partial D$ is called regular accessible (RA) in $D\in\{D_{\lambda}\}$ if there exists a regular $\omega_{\cdot}^{\widetilde{D}}\in\{\omega_{\cdot}^{D_{\lambda}}\}$ with $\widetilde{D}\subseteq D$ such that $y\in\partial\widetilde{D}\cap\partial D$.
(3) A harmonic measure system $\{\omega_{\cdot}^{D_{\lambda}}\}$ on $X$ is called regular closed if for any $D\in\{D_{\lambda}\}$, there exists an increasing regular sequence $\{D_{k}\}$ such that $D=\mathop{\cup}\limits_{k=1}^{\infty}D_{k}$.
In a harmonic system, the following proposition holds.
Proposition 2.8.
Suppose in a harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$, $D_{1},D_{2},D_{1}\cup D_{2}\in\{D_{\lambda}\}$. If $D_{1}$ and $D_{2}$ are regular, then $D_{1}\cup D_{2}$ is regular.
Proof.
Since $D_{1}$ and $D_{2}$ are regular, for any $f_{1}\in L(\partial D_{1};\omega_{\cdot}^{D_{1}})$ continuous at $y_{1}\in\partial D_{1}$, $f_{2}\in L(\partial D_{2};\omega_{\cdot}^{D_{2}})$ continuous at $y_{2}\in\partial D_{2}$,
$\lim_{D_{1}\ni x_{n}\rightarrow y_{1}}\omega_{x_{n}}^{D_{1}}(f_{1})=f_{1}(y_{1}),\quad\lim_{D_{2}\ni x_{n}\rightarrow y_{2}}\omega_{x_{n}}^{D_{2}}(f_{2})=f_{2}(y_{2}).$
If $D_{1}\cap D_{2}=\emptyset$ the conclusion holds obviously.
If $D_{1}\cap D_{2}\not=\emptyset$, then for any $y\in\partial(D_{1}\cup D_{2})$, $y\in\partial D_{1}\backslash\partial D_{2}$ or $y\in\partial D_{2}\backslash\partial D_{1}$ or $y\in\partial D_{1}\cap\partial D_{2}$.
If $y\in\partial D_{1}\backslash\partial D_{2}$, taking the above $f_{1}(\cdot):=\omega_{\cdot}^{D_{1}\cup D_{2}}(f)$ on $\partial D_{1}$ for any $f\in L(\partial(D_{1}\cup D_{2});\omega_{\cdot}^{D_{1}\cup D_{2}})$ continuous at $y\in\partial D_{1}\backslash\partial D_{2}$, then $f_{1}\in L(\partial D_{1};\omega_{\cdot}^{\partial D_{1}})$ is continuous at $y\in\partial D_{1}$. Then by the compatibility of $\{\omega_{\cdot}^{D}\}$, we get
$\lim_{D_{1}\cup D_{2}\ni x_{n}\rightarrow y\in\partial D_{1}\backslash\partial D_{2}}\omega_{x_{n}}^{D_{1}\cup D_{2}}(f)=\lim_{D_{1}\ni x_{n}\rightarrow y\in\partial D_{1}}\omega_{x_{n}}^{D_{1}}(f_{1})=f_{1}(y)=f(y),$
which means $y\in\partial D_{1}\backslash\partial D_{2}$ is regular relative to $D_{1}\cup D_{2}$.
If $y\in\partial D_{2}\backslash\partial D_{1}$, similar argument of the above also applies.
If $y\in\partial D_{1}\cap\partial D_{2}$, denoting $\{x_{n_{1}}\}=\{x_{n}\}\cap(D_{1}\backslash\bar{D}_{2})$, $\{x_{n_{2}}\}=\{x_{n}\}\cap(D_{2}\backslash\bar{D}_{1})$ and $\{x_{n_{3}}\}=\{x_{n}\}\cap(\bar{D}_{1}\cap\bar{D}_{2})$, then for any $f\in L(\partial(D_{1}\cup D_{2});\omega_{\cdot}^{D_{1}\cup D_{2}})$ continuous at $y\in\partial D_{1}\cap\partial D_{2}$, similar argument yields
$\lim_{(D_{1}\backslash D_{2})\ni x_{n_{1}}\rightarrow y\in\partial D_{1}\cap\partial D_{2}}\omega_{x_{n_{1}}}^{D_{1}\cup D_{2}}(f)=\lim_{D_{1}\ni x_{n_{1}}\rightarrow y\in\partial D_{1}}\omega_{x_{n_{1}}}^{D_{1}}(f_{1})=f_{1}(y)=f(y),$
$\lim_{(D_{2}\backslash D_{1})\ni x_{n_{2}}\rightarrow y\in\partial D_{2}\cap\partial D_{1}}\omega_{x_{n_{2}}}^{D_{1}\cup D_{2}}(f)=\lim_{D_{2}\ni x_{n_{2}}\rightarrow y\in\partial D_{2}}\omega_{x_{n_{2}}}^{D_{2}}(f_{2})=f_{2}(y)=f(y),$
$\lim_{(D_{1}\cap D_{2})\ni x_{n_{3}}\rightarrow y\in\partial D_{1}\cap\partial D_{2}}\omega_{x_{n_{3}}}^{D_{1}\cup D_{2}}(f)=\lim_{D_{1}\ni x_{n_{3}}\rightarrow y\in\partial D_{1}}\omega_{x_{n_{3}}}^{D_{1}}(f_{1})=f_{1}(y)=f(y).$
The above three equations imply
$\lim_{D_{1}\cup D_{2}\ni x_{n}\rightarrow y\in\partial D_{1}\cap\partial D_{2}}\omega_{x_{n}}^{D_{1}\cup D_{2}}(f)=f(y),$
which means $y\in\partial D_{1}\cap\partial D_{2}$ is regular relative to $D_{1}\cup D_{2}$. ∎
Remark 2.2.
If $D_{1}$ and $D_{2}$ are regular in a harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$, we can not get $D_{1}\cap D_{2}$ is regular in general provided $D_{1}\cap D_{2}\in\{D_{\lambda}\}$.
Theorem 2.9.
Suppose the following two conditions hold: (1) The harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ in $\mathbb{R}^{n}$ is translation closed and invariant, that is, for any $x_{0}\in\mathbb{R}^{n}$ and $D\in\{D_{\lambda}\}$, such that $D+\{x_{0}\}\in\{D_{\lambda}\}$; for any $x\in D$ and any $A\subseteq\partial D$, such that $\omega_{x+x_{0}}^{D+\{x_{0}\}}(A+\{x_{0}\})=\omega_{x}^{D}(A)$; (2) $\{D_{\lambda}\}$ is closed about the finite operation $\cup$”; (3) There exists a $\{\omega_{\cdot}^{O_{k}}\}$-regular open set sequence $\{O_{k}\}\subseteq\{D_{\lambda}\}$ in $\mathbb{R}^{n}$ such that $\lim\limits_{k\rightarrow\infty}\mathrm{diam}\,O_{k}=0$. Then $\{\omega_{\cdot}^{D_{\lambda}}\}$ is regular closed.
Proof.
Suppose $D\in\{D_{\lambda}\}$. Let $\{K_{l}\}$ be an increasing sequence of compact sets such that $D=\mathop{\cup}\limits_{l}K_{l}$. Then there exists a finite covering $\{O^{l}_{i}\}_{i=1}^{k_{l}}$ of $K_{l}$, where $O^{l}_{i}\in\{D_{\lambda}\}$, $i=1,2,\cdots,k_{l}$ are regular and $\bar{O}^{l}_{i}\subset D$. By Proposition 2.8 $D_{l}:=\mathop{\cup}\limits_{j=1}^{k_{l}}O^{l}_{j}\in\{D_{\lambda}\}$ is regular. It is clear that $\bar{D}_{l}\subset D$ and $D=\mathop{\cup}\limits_{l=1}^{\infty}D_{l}$. ∎
Theorem 2.10 (Strong Maximum Principle for open set).
Suppose the harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ is regular closed, $D\in\{D_{\lambda}\}$ and $f\in L(\partial D,\omega_{\cdot}^{D})\cap L^{\infty}(\partial D,\omega_{\cdot}^{D})$. If there exists an $x_{0}\in D$ such that
$\omega^{D}_{x_{0}}(f)=\sup_{x\in D}\omega_{x}^{D}(f),$
then $\omega_{x}^{D}(f)$ is constant for $x\in D$.
Proof.
Since the harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ is regular closed, there exists an increasing regular open set sequence $\{D_{k}\}$ such that $D=\mathop{\cup}\limits_{k=1}^{\infty}D_{k}$. By $x_{0}\in D$, we can find a $\widetilde{K}\in\mathbb{N}$, for any $k\geq\widetilde{K}$ such that $x_{0}\in D_{k}$.
For any $x\in D$, we prove $\omega_{x}^{D}(f)=\omega_{x_{0}}^{D}(f)$. In fact, since $x\in D$, there exists a $K\geq\widetilde{K}$ such that $x_{0},x\in D_{K}$. Then by the compatibility of the harmonic system $\{\omega_{\cdot}^{D}\}$, we get
$\omega_{x_{0}}^{D_{K}}(\omega_{\cdot}^{D}(f))=\omega_{x_{0}}^{D}(f)=\sup_{x\in D}\omega_{x}^{D}(f)\mathop{=}\limits^{x_{0}\in D_{K}}\sup_{x\in D_{K}}\omega_{x}^{D}(f)=\sup_{x\in D_{K}}\omega_{x}^{D_{K}}(\omega_{\cdot}^{D}(f)).$
Since $D_{K}$ is regular and $\omega_{\cdot}^{D}(f)$ is continuous on $\partial D_{K}$, by Theorem 2.7, we get
$\omega_{x}^{D}(f)=\omega_{x}^{D_{K}}(\omega_{\cdot}^{D}(f))=\omega_{x_{0}}^{D_{K}}(\omega_{\cdot}^{D}(f))=\omega_{x_{0}}^{D}(f),$
which completes the proof. ∎
From the proof of the above theorem, we can get the following corollary.
Corollary 2.1.
Suppose $\{\omega_{\cdot}^{D_{\lambda}}\}$ is a harmonic system and $D\in\{D_{\lambda}\}$. If there exists an increasing regular open set sequence $\{D_{k}\}$ such that $D=\mathop{\cup}\limits_{k=1}^{\infty}D_{k}$, $f\in L(\partial D,\omega_{\cdot}^{D})\cap L^{\infty}(\partial D,\omega_{\cdot}^{D})$. If there exists an $x_{0}\in D$ such that
$\omega^{D}_{x_{0}}(f)=\sup_{x\in D}\omega_{x}^{D}(f),$
then $\omega_{x}^{D}(f)$ is constant for $x\in D$.
2.4. Harnack Domains
In this subsection we consider the domains in which the Harnack’s inequality holds.
Definition 2.4.
Given a harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ on the Hausdorff space $X$. Suppose $D\in\{D_{\lambda}\}$ with an subset $\widetilde{D}\subseteq D$. $H:\widetilde{D}\subseteq D\mapsto H(\widetilde{D};D)\in[1,+\infty]$ defined by
$H(\widetilde{D};D):=\sup_{x,y\in\widetilde{D}\atop 0\leq f\in L(\partial D,\omega_{\cdot}^{D})\backslash\{0\}}\frac{\omega_{x}^{D}(f)}{\omega_{y}^{D}(f)}$
is called the Harnack index of the subset $\widetilde{D}$ relative to $D$. Denote
$D_{\alpha}:=\{\widetilde{D}\subseteq D:\,H(\widetilde{D};D)=\alpha\}.$
The elements of $D_{\alpha}$ are called the Harnack subsets with index $\alpha$ relative to $D$.
By the definition of Harnack index, the following propositions holds.
Proposition 2.11.
  1. (1)
    For any $D\in\{D_{\lambda}\}$, $H(D;D)=+\infty$;
  2. (2)
    If $D_{1},D_{2}\in\{D_{\lambda}\}$ with a subset $D_{3}\subseteq D_{2}\subseteq D_{1}$, then $H(D_{3};D_{1})\leq H(D_{3};D_{2})$ and $H(D_{3};D_{1})\leq H(D_{2};D_{1})$;
  3. (3)
    For any $D\in\{D_{\lambda}\}$, any $f\in L(\partial D,\omega_{\cdot}^{D})\backslash\{0\}$ with $f\geq 0$ and any $x,y\in\widetilde{D}\subseteq D$,
    $\omega_{x}^{D}(f)\leq H(\widetilde{D};D)\omega_{y}^{D}(f).$
By $\omega^{D}_{x}\ll\omega^{D}_{y}$ in the harmonic measure system, Radon-Nikodym Theorem implies $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$ exists. The boundedness of $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$ implies $H(\widetilde{D};D)<+\infty$.
Theorem 2.12 (General Harnack’s inequality).
Suppose $D\in\{D_{\lambda}\}$. If $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$ is bounded for any $x,y\in\widetilde{D}\subseteq D$, then $H(\widetilde{D};D)<+\infty$.
Proof.
By the boundedness of $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$, for any $f\in L(\partial D,\omega_{\cdot}^{D})\backslash\{0\}$ with $f\geq 0$ and any $x,y\in\widetilde{D}\subseteq D$, Radon-Nikodym Theorem implies
$\frac{\omega^{D}_{x}(f)}{\omega^{D}_{y}(f)}=\frac{\int f\,\mathrm{d}\omega_{x}^{D}}{\int f\,\mathrm{d}\omega_{y}^{D}}=\frac{\int f\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}\,\mathrm{d}\omega_{y}^{D}}{\int f\,\mathrm{d}\omega_{y}^{D}}\leq\frac{\int Mf\,\mathrm{d}\omega_{y}^{D}}{\int f\,\mathrm{d}\omega_{y}^{D}}=M.$
Then $H(\widetilde{D};D)\leq M<\infty$. ∎
Corollary 2.2 (Harnack’s inequality).
Suppose $D\in\{D_{\lambda}\}$. If $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$ is continuous for any $x,y\in D$. Then $H(\widetilde{D};D)<+\infty$ for any subset $\widetilde{D}\subset\subset D$.
In the following of this subsection, we limit our discussion in $X=\mathbb{R}^{n}$. We make the following assumption:
  1. (TSCI)
    (Translation and Scaling Closed & Invariant) The harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ in $\mathbb{R}^{n}$ is closed and invariant under the transformation of translation and scaling, that is, for any $x_{0},\nu_{0}\in\mathbb{R}^{n}$ and $k\in\mathbb{R}(k\not=0)$, define the translation and scaling map for any $x\in\mathbb{R}^{n}$ by $\nu_{\nu_{0},x_{0},k}(x):=\nu_{0}+k(x-x_{0})$, then
    1. (1)
      (Translation and Scaling Closed) For any open $D\in\{D_{\lambda}\}$, $\nu_{\nu_{0},x_{0},k}(D)\in\{D_{\lambda}\}$;
    2. (2)
      (Translation and Scaling Invariant) For any $\omega_{\cdot}^{D}$-measurable set $A\subseteq\partial D$, $\nu_{\nu_{0},x_{0},k}(A)$ is $\omega_{\cdot}^{\nu_{\nu_{0},x_{0},k}(D)}$-measurable with
      $\omega_{\nu_{\nu_{0},x_{0},k}(x)}^{\nu_{\nu_{0},x_{0},k}(D)}(\nu_{\nu_{0},x_{0},k}(A))=\omega_{x}^{D}(A).$
Roughly speaking, the TSCI assumption on the harmonic measure system means that the harmonic measure $\omega_{\cdot}^{D}$ in the harmonic system depends on the shape of $D\in\{D_{\lambda}\}$ but independent on its size and location in $\mathbb{R}^{n}$.
Theorem 2.13 (Harnack’s Principle in $\mathbb{R}^{n}$).
Suppose the harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ in $\mathbb{R}^{n}$ satisfies TSCI and $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$ is continuous for any $x,y\in D\in\{D_{\lambda}\}$. If $D,D_{1}\in\{D_{\lambda}\}$ with a subset $D_{2}\subset\subset D_{1}\subseteq D$, then $H(\nu(D_{2});D)\leq H(D_{2};D_{1})<+\infty$ provided $\nu(D_{1})\subseteq D$ for any translation and scaling map $\nu$. Equivalently, for any $f\in L(\partial D;\omega_{\cdot}^{D})$ and any $x,y\in D_{2}$, there exists a constant $C:=H(D_{2};D_{1})$ such that
$\omega_{x}^{D}(f)\leq C\omega_{y}^{D}(f).$
Proof.
The TSCI assumption on the harmonic system $\{\omega_{\cdot}^{D}\}$, Proposition 2.11 (2) and Theorem 2.12 imply
$H(\nu(D_{2});D)\leq H(\nu(D_{2});\nu(D_{1}))=H(D_{2};D_{1})<+\infty.$
Remark 2.3.
In the applications when we use the Harnack Principle in $\mathbb{R}^{n}$, we always say Harnack inequality holds for the constant $C$ depending on the shape of $D_{2}\subset\subset D_{1}$. See a special inner cone case of the Harnack principle in Proposition 4.2.
3. Dirichlet boundary problem in graph theory
Let $X$ be a finite set, whose elements are called vertexes. Define a weight function $\mu:X\times X\rightarrow[0,\infty),(x,y)\mapsto\mu_{xy}$ which is symmetric, has zero diagonal. $\mu_{xy}>0$ represents there is an edge between $x$ and $y$ and $\mu_{xy}=0$ represents there is no edge between $x$ and $y$. $X$ with the edges induced by $\mu$ is a edge weighted graph, denoted by $(X,\mu)$. If $\mu_{xy}>0$, we write $x\sim y$ and let $xy$ and $yx$ be the oriented edges of the graph. We assume $(X,\mu)$ is connected, that is for any two vertexes $x,y\in X$, there is a path $x=x_{0}\sim x_{1}\sim x_{2}\sim\cdots\sim x_{n}=y$.
It is well known that $(X,\mu)$ can be embedded on a compact surface in an Euclidean space. We define the topology of $(X,\mu)$ be the induced topology by the embedded surface. For any connected subgraph $(X^{\prime},\mu^{\prime}:=\mu|_{X^{\prime}\times X^{\prime}})$ of $(X,\mu)$, denote $\bar{X}^{\prime}:=X^{\prime}\cup B$, where $B:=\{y\in X\backslash X^{\prime}:y\sim x\in X^{\prime}\}$ is called the relative boundary of $(X^{\prime},\mu^{\prime})$. $(\bar{X}^{\prime},\bar{\mu}^{\prime}:=\mu|_{\bar{X}^{\prime}\times\bar{X}^{\prime}})^{\circ}:=(\bar{X}^{\prime},\bar{\mu}^{\prime})\backslash(B,\mu_{B}:=\mu|_{B\times B})$ is called open subgraph of $(X,\mu)$ containing $(X^{\prime},\mu^{\prime})$. Under the topology induced by the topology of $(X,\mu)$, $\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})^{\circ}=B$.
Suppose $(X^{\prime},\mu^{\prime})$ is a connected subgraph of $(X,\mu)$. We define the Laplacian $\Delta_{(\bar{X}^{\prime},\bar{\mu}^{\prime})}$ on the set $U(\bar{X}^{\prime}):=\{u:\bar{X}^{\prime}\rightarrow\mathbb{R}\}$ for any $x\in(\bar{X}^{\prime},\bar{\mu}^{\prime})^{\circ}\cap\bar{X}^{\prime}=X^{\prime}$ by
$\Delta_{(\bar{X}^{\prime},\bar{\mu}^{\prime})}u(x)=\sum_{x\sim y\in\bar{X}^{\prime}}\bar{\mu}^{\prime}(x,y)(u(x)-u(y)).$
A function $u:\bar{X}^{\prime}\rightarrow\mathbb{R}$ is called harmonic at $x\in X^{\prime}$ if $u\in U(\bar{X}^{\prime})$ satisfies $\Delta_{(\bar{X}^{\prime},\bar{\mu}^{\prime})}u(x)=0$.
The Dirichlet boundary problem on $(\bar{X}^{\prime},\bar{\mu}^{\prime})$ is, for a given $f:B\rightarrow\mathbb{R}$, to find a function $u\in U(\bar{X}^{\prime})$ such that
$(P)\left\{\begin{aligned} &\Delta_{(\bar{X}^{\prime},\bar{\mu}^{\prime})}u(x)=0&&\text{ for }x\in X^{\prime};\\ &u(x)=f(x)&&\text{ for }x\in B.\end{aligned}\right.$
Lemma 3.1.
For any $x\in(\bar{X}^{\prime},\bar{\mu}^{\prime})^{\circ}\cap\bar{X}^{\prime}$, the solution $u$ of $(P)$ satisfies
$u(x)=\sum_{x\sim y\in\bar{X}^{\prime}}\frac{\bar{\mu}^{\prime}(x,y)}{\sum\limits_{x\sim z\in\bar{X}^{\prime}}\bar{\mu}^{\prime}(x,z)}u(y)$
Theorem 3.1.
If $(\bar{X}^{\prime},\bar{\mu}^{\prime})$ is connected and $\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})\not=\emptyset$, then the solution to $(P)$ exists uniquely.
Proof.
Denote $(\bar{X}^{\prime},\bar{\mu}^{\prime})^{\circ}\cap\bar{X}^{\prime}=\{x_{i}\}_{i=1}^{n}$, $\bar{\mu}^{\prime}(x_{i},x_{j}):=a_{ij}$. By Lemma 3.1, the existence and uniqueness of the solution to $(P)$ is equivalent to the existence and uniqueness of the solution to linear system $Ax=y$, where
$A:=\begin{bmatrix}k_{1}+\sum_{j=1}^{n}a_{1j}&-a_{12}&-a_{13}&\cdots&-a_{1n}\\ -a_{21}&k_{2}+\sum_{j=1}^{n}a_{2j}&-a_{23}&\cdots&-a_{2n}\\ -a_{31}&-a_{32}&k_{3}+\sum_{j=1}^{n}a_{3j}&\cdots&-a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -a_{n1}&-a_{n2}&-a_{n3}&\cdots&k_{n}+\sum_{j=1}^{n}a_{nj}\end{bmatrix},$
$k_{i}:=\sum_{x_{j}\sim y\in\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})}\bar{\mu}^{\prime}(x_{i},y),$
$x:=\begin{bmatrix}u(x_{1})\\ u(x_{2})\\ \vdots\\ u(x_{n})\end{bmatrix},$
$y:=\begin{bmatrix}\sum\limits_{x_{1}\sim y\in\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})}\bar{\mu}^{\prime}(x_{1},y)f(y)\\ \sum\limits_{x_{2}\sim y\in\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})}\bar{\mu}^{\prime}(x_{2},y)f(y)\\ \vdots\\ \sum\limits_{x_{n}\sim y\in\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})}\bar{\mu}^{\prime}(x_{n},y)f(y)\end{bmatrix}.$
It is clear that $k_{i}\geq 0$, $a_{ij}\geq 0$.
We will prove $A$ is positive, that is, for any $0\not=\xi=(\xi_{1},\xi_{2},\cdots,\xi_{n})\in\mathbb{R}^{n}$, $\xi A\xi^{T}>0$. In fact,
$\xi A\xi^{T}=\sum_{i=1}^{n}k_{i}\xi_{i}^{2}+\sum_{i,j=1}^{n}a_{ij}(\xi_{i}-\xi_{j})^{2}.$
(1) If $\sum\limits_{i,j=1}^{n}a_{ij}(\xi_{i}-\xi_{j})^{2}>0$, then $A$ is positive.
(2) if $\sum\limits_{i,j=1}^{n}a_{ij}(\xi_{i}-\xi_{j})^{2}=0$, then the connectivity of $(\bar{X}^{\prime},\bar{\mu}^{\prime})$ implies that $\xi_{1}=\xi_{2}=\cdots=\xi_{n}\not=0$. By $\partial(\bar{X}^{\prime},\bar{\mu}^{\prime})\not=\emptyset$, we know there exists a $k_{l}\not=0$. Then $\sum\limits_{i=1}^{n}k_{i}\xi_{i}^{2}>0$, which implies that $A$ is positive.
By (1) and (2), $A$ is positive. Then $|A|\not=0$, which implies the existence and uniqueness of the solution to $Ax=y$, equivalently, solution to $(P)$ exists uniquely. ∎
For a simple graph $(X,\mu)$ with $X=\{x_{1},x_{2}\}$ and $\mu(x_{1},x_{2})>0$, we add a new vertex $x_{\lambda}\in x_{1}x_{2}$ in $X$ in the following way:
(3.1) $\frac{\widetilde{\mu}(x_{1},x_{\lambda})}{\widetilde{\mu}(x_{\lambda},x_{2})}=\frac{1-\lambda}{\lambda}:=\frac{|x_{\lambda}x_{2}|}{|x_{1}x_{\lambda}|}.$
(3.2) $\frac{1}{\widetilde{\mu}(x_{1},x_{\lambda})}+\frac{1}{\widetilde{\mu}(x_{\lambda},x_{2})}=\frac{1}{\mu(x_{1},x_{2})},$
Then $(\widetilde{X}:=\{x_{1},x_{\lambda},x_{2}\},\widetilde{\mu})$ is a new edge weighted graph. (3.1) ensures the following lemma holds.
Lemma 3.2.
In the graph $(\widetilde{X},\widetilde{\mu})$, $u(x_{\lambda}):=(1-\lambda)u(x_{1})+\lambda u(x_{2})$ is harmonic at $x_{\lambda}$, where $\lambda$ is defined in (3.1).
Proof.
By the Lemma 3.1 and (3.1), if $\Delta_{(\widetilde{X},\widetilde{\mu})}u(x_{\lambda})=0$, then
$u(x_{\lambda})$ $=\frac{\widetilde{\mu}(x_{\lambda},x_{1})}{\widetilde{\mu}(x_{\lambda},x_{1})+\widetilde{\mu}(x_{\lambda},x_{2})}u(x_{1})+\frac{\widetilde{\mu}(x_{\lambda},x_{2})}{\widetilde{\mu}(x_{\lambda},x_{1})+\widetilde{\mu}(x_{\lambda},x_{2})}u(x_{2})$
$=(1-\lambda)u(x_{1})+\lambda u(x_{2}).$
(3.1) and (3.2) ensures the following lemma holds.
Lemma 3.3.
Suppose $x_{1}x_{2}$ is an edge in the subgraph $(X^{\prime},\mu^{\prime})$ of $(X,\mu)$, $x_{1}\in(\bar{X}^{\prime},\bar{\mu}^{\prime})^{\circ}$ and $u$ is harmonic at $x_{1}$. If $x_{\lambda}\in x_{1}x_{2}$ and $u(x_{\lambda}):=(1-\lambda)u(x_{1})+\lambda u(x_{2})$, then $u$ is harmonic at $x_{1}$ in $(\overline{X^{\prime}\cup\{x_{\lambda}\}},\widetilde{\bar{\mu}^{\prime}})$, where $\widetilde{\bar{\mu}^{\prime}}(x_{1},x_{\lambda})$ and $\widetilde{\bar{\mu}^{\prime}}(x_{2},x_{\lambda})$ satisfies conditions as (3.1) and (3.2), on the other vertexes $\widetilde{\bar{\mu}^{\prime}}=\bar{\mu}^{\prime}$.
Proof.
Since $u$ is harmonic at $x_{1}$ in $(\bar{X}^{\prime},\bar{\mu}^{\prime})$,
00 0 $=\Delta_{(\bar{X}^{\prime},\bar{\mu}^{\prime})}u(x_{1})=\sum_{x_{1}\sim y\in\bar{X}^{\prime}}\bar{\mu}^{\prime}(x_{1},y)(u(x_{1})-u(y))$
=$\bar{\mu}^{\prime}(x_{1},x_{2})(u(x_{1})-u(x_{2}))+\sum_{x_{1}\sim y\not=x_{2},y\in\bar{X}^{\prime}}\bar{\mu}^{\prime}(x_{1},y)(u(x_{1})-u(y)).$
In $(\overline{X^{\prime}\cup\{x_{\lambda}\}},\widetilde{\bar{\mu}^{\prime}})$, we get
$\Delta_{(\overline{X^{\prime}\cup\{x_{\lambda}\}},\widetilde{\bar{\mu}^{\prime}})}u(x_{1})$ $=\sum_{x_{1}\sim y\in\overline{X^{\prime}\cup\{x_{\lambda}\}}}\widetilde{\bar{\mu}^{\prime}}(x_{1},y)(u(x_{1})-u(y))$
$=\widetilde{\bar{\mu}^{\prime}}(x_{1},x_{\lambda})(u(x_{1})-u(x_{\lambda}))$
$\quad\quad+\sum_{x_{1}\sim y\not=x_{\lambda},y\in\overline{X^{\prime}\cup\{x_{\lambda}\}}}\widetilde{\bar{\mu}^{\prime}}(x_{1},y)(u(x_{1})-u(y))$
$=A(x_{1},x_{\lambda},x_{2})\bar{\mu}^{\prime}(x_{1},x_{2})(u(x_{1})-u(x_{2}))$
$\quad\quad+\sum_{x_{1}\sim y\not=x_{2},y\in\bar{X}^{\prime}}\bar{\mu}^{\prime}(x_{1},y)(u(x_{1})-u(y)),$
where
$A(x_{1},x_{\lambda},x_{2}):=\frac{\widetilde{\bar{\mu}^{\prime}}(x_{1},x_{\lambda})}{\bar{\mu}^{\prime}(x_{1},x_{2})}\frac{u(x_{1})-u(x_{\lambda})}{u(x_{1})-u(x_{2})}.$
By (3.1) and (3.2), it is easy to get $A(x_{1},x_{\lambda},x_{2})=\dfrac{1}{\lambda}\cdot\lambda=1$. Then $\Delta_{(\overline{X^{\prime}\cup\{x_{\lambda}\}},\widetilde{\bar{\mu}^{\prime}})}u(x_{1})=\Delta_{(\bar{X}^{\prime},\bar{\mu}^{\prime})}u(x_{1})=0$, which completes the proof. ∎
In general, we can extend the harmonic function on vertexes $X^{\prime}$ to points on edges of $(\bar{X}^{\prime},\bar{\mu}^{\prime})^{\circ}$ by Lemma 3.2 and Lemma 3.3. Then for any open connected set $O$ of the compact surface the graph $(X,\mu)$ embedded on, and any $x\in D:=(X,\mu)\cap O$ with $D$ being connected and $\partial D\not=\emptyset$, given an $f$ on $\partial D$, there exists a continuous function $\widetilde{u}(x)$ on $D$, such that $\Delta_{((X\cap O)\cup\{x\},\widetilde{\mu})}\widetilde{u}(y)=0$ for $y\in(X\cap O)\cup\{x\}$, where $\widetilde{\mu}$ is induced by Lemma 3.3. Denote $\omega^{D}_{x}(f):=\widetilde{u}(x)$. Then the topological space $((X,\mu),\tau)$ is a Hausdorff space, where $\tau$ is induced by the topology on the compact surface the graph $(X,\mu)$ embedded on. By Theorem 3.1, Lemma 3.2 and Lemma 3.3 we get the following theorem.
Theorem 3.2.
(1) $\{\omega_{\cdot}^{D}\}$ is a harmonic system on $(X,\mu)$; (2) Any open set in $(X,\mu)$ is regular; (3) The harmonic system $\{\omega_{\cdot}^{D}\}$ is regular closed; (4) Strong maximum principle holds for any connected open set with nonempty boundary in $(X,\mu)$.
4. Dahlber’s theory
In this section, we generalize Dahlberg’s theory [2] under the axiomatized harmonic measure theory. The proof is based on [2].
4.1. Main theorem
Suppose $\{\omega_{\cdot}^{D_{\lambda}}\}$ is a harmonic system on $\mathbb{R}^{n}$. In this section the harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ satisfies the following assumptions (H):
  1. (H1)
    $\{D_{\lambda}\}=\{\text{All the bounded open set in }\mathbb{R}^{n}\}$.
  2. (H2)
    Any bounded Lipschitz domain $D\subseteq\mathbb{R}^{n}$ is regular relative to $\omega_{\cdot}^{D}$.
  3. (H3)
    The harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ is translation and scaling closed & invariant (TSCI), see the assumption TSCI in Section 2.4.
  4. (H4)
    For any bounded open $D\subseteq\mathbb{R}^{n}$, $\frac{\mathrm{d}\omega^{D}_{x}}{\mathrm{d}\omega^{D}_{y}}$ is continuous for any $x,y\in D$.
Remark 4.1.
(a) Under the assumptions (H2) and (H3), by Theorem 2.9, $\mathbb{R}^{n}$ is regular closed. By Theorem 2.10 Strong Maximal Principle is available. (b) By the discussion of Section 2.4, under the assumption (H4) Harnack inequality is available.
A Lipschitz domain $D$ is called starlike about $P_{0}$ if each ray emanating from $P_{0}$ intersects $\partial D$ exactly once and the local coordinate system can be taken with its $y$-axis along $QP_{0}$ for any $Q\in\partial D$.
Definition 4.1.
($B_{1+\delta}$ condition) For $\delta>0$, we say that $\mu_{2}$ satisfies $B_{1+\delta}(\mu_{1})$ condition relative to $\mu_{1}$ denoted by $\mu_{2}\in B_{1+\delta}(\mu_{1})$ if $\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\in L^{\delta}(\partial D,\mu_{2})$ exists and there exits a constant $C>0$ such that
$\fint_{Q}\left(\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)^{\delta}\,\mathrm{d}\mu_{2}\leq C\left(\frac{\mu_{2}(Q)}{\mu_{1}(Q)}\right)^{\delta}$
for any $Q:=B\cap\partial D\not=\emptyset$.
The main result of this section is the following theorem.
Theorem 4.1.
Suppose the harmonic system $\{\omega_{\cdot}^{D_{\lambda}}\}$ satisfies (H), $D\subset\mathbb{R}^{n}\,(n\geq 3)$ is a bounded Lipschitz domain, starlike about the origin $O\in D$. If $\mu$ is a positive measure on $D$ and there exists a $\delta>0$ such that for any $x\in D$, $\omega_{x}^{D}\in B_{1+\delta}(\omega_{O}^{D})$, $(1+\delta)^{\prime}<p<\infty$, then the following conditions are equivalent:
  1. (i)
    $\mu$ is a Carleson measure in $D$, that is, $\exists\,M>0,\forall\,P\in\partial D,\forall\,r>0$ such that
    $\mu(B_{P,r}\cap D)\leq M\omega_{O}(B_{P,r}\cap\partial D);$
  2. (ii)
    $H:f\mapsto\omega_{\cdot}^{D}(f)$ is strongly $(p,p)$, that is, $\exists\,K>0,\forall\,f\in L^{p}(\omega_{O}^{D})$ such that
    $\int_{D}|(Hf)(x)|^{p}\,\mathrm{d}\mu(x)\leq K\int_{\partial D}|f|^{p}\,\mathrm{d}\omega_{O}^{D}.$
4.2. The proof of (ii) $\Rightarrow$ (i)
Lemma 4.1.
Suppose $D$ is a Lipschitz domain. Then there exists a positive constant $\delta=\delta(D)\leq\frac{1}{2}$ such that
$\omega_{x}^{D}(B_{P,r}\cap\partial D)\geq\delta$
for any $P\in\partial D$, any $r>0$ and any $x\in B_{P,\delta r}\cap D$.
Proof.
Let $\Delta^{\prime}$ be the part of $\partial B_{P,r}$ outside $D$ and $\Delta_{P,r}:=B_{P,r}\cap\partial D$. Then $\omega_{x}^{D}(\Delta_{P,r})\geq\omega_{x}^{B_{P,r}}(\Delta^{\prime})$ for all $x\in\partial(D\cap B_{P,r})$. By the Strong Maximum Principle 2.10, we get
$\omega_{x}^{D}(\Delta_{P,r})\geq\omega_{x}^{B_{P,r}}(\Delta^{\prime}),$
for any $x\in B_{P,r}\cap D$. Take $\delta_{0}:=\inf\limits_{x\in B_{P,\frac{1}{2}r}\cap D}\omega_{x}^{B_{P,r}}(\Delta^{\prime}\cap\Gamma_{D})$, where $\Gamma_{D}$ is the cone outside D with its vertex at $P$ satisfying $\Delta^{\prime}\cap\Gamma_{D}=\partial B_{P,r}\cap\Gamma_{D}$ (since $\partial D$ is Lipschitz, such cone exists). By the Strong Maximum Principle 2.10, it is clear that $\delta_{0}<1$. And $\delta_{0}$ depends only on $D$ by the translation and scaling invariant property of the harmonic system $\{\omega_{\cdot}^{D}\}$ and for any $x\in B_{P,\frac{1}{2}r}\cap D$,
$\omega_{x}^{D}(\Delta_{P,r})\geq\omega_{x}^{B_{P,r}}(\Delta^{\prime})\geq\delta_{0}.$
Take $\delta:=\min\{\frac{1}{2},\delta_{0}\}$. Then for any $x\in B_{P,\delta r}\cap D$,
$\omega_{x}^{D}(\Delta_{P,r})\geq\delta,$
where $\delta$ depends only on $D$. ∎
Since $D$ is a bounded Lipschitz domain, starlike about the origin $O\in D$. For any $P\in\partial D$, there exists an open cycle cylinder centered at $P$ with small radius $r>0$, height $sr>0$ and axis along line $OP$, denoted by $C_{P,r}=C_{P,r,sr}$. The parameter $s$ depends on the the Lipschitz constant $L$ of $\partial D$ such that cylinder $C_{P,r}$ satisfies one bottom is contained in $D$ and the other is outside. For example, we can take $s=2L$. Denote $\Delta_{P,r}:=C_{P,r}\cap\partial D$. The center point of the bottom of $C_{P,r}$ inside $D$ is denoted by $T_{P,r}$.
We can get another form of Lemma 4.1 as follows.
Lemma 4.2.
Suppose $D$ is a Lipschitz domain. Then there exists a positive constant $\delta=\delta(s,D)\leq\frac{1}{2}$ such that
$\omega_{x}^{D}(\Delta_{P,r})\geq\delta$
for any $x\in C_{P,\delta r}\cap D$.
Proof.
Let $\Delta^{\prime}$ be the bottom of $C_{P,r}$ outside $D$. Then $\chi_{\Delta_{P,r}}(x)\geq\omega_{x}^{C_{P,r}}(\Delta^{\prime})$ for all $x\in\partial(D\cap C_{P,r})$. By the Strong Maximum Principle 2.10, we get
$\omega_{x}^{D}(\Delta_{P,r})\geq\omega_{x}^{C_{P,r}}(\Delta^{\prime}),$
for any $x\in C_{P,r}\cap D$. Take $\delta_{0}:=\inf\limits_{x\in C_{P,\frac{1}{2}r}\cap D}\omega_{x}^{C_{P,r}}(\Delta^{\prime})$. By the Strong Maximum Principle 2.10, it is clear that $\delta_{0}<1$, and $\delta_{0}$ depends only on $s$ and $D$ by the translation and scaling invariant property of the harmonic system $\{\omega_{\cdot}^{D}\}$ and for any $x\in C_{P,\frac{1}{2}r}\cap D$,
$\omega_{x}^{D}(\Delta_{P,r})\geq\omega_{x}^{C_{P,r}}(\Delta^{\prime})\geq\delta_{0}.$
Take $\delta:=\min\{\frac{1}{2},\delta_{0}\}$. Then for any $x\in C_{P,\delta r}\cap D$,
$\omega_{x}^{D}(\Delta_{P,r})\geq\delta,$
where $\delta$ depends only on $s$ and $D$. ∎
The Proof of (ii) $\Rightarrow$ (i).
For any $p\geq 2,P\in\partial D,r>0$ and $x\in B_{P,r}\cap D$, $\omega^{D}_{x}(B_{P,\frac{r}{\delta}}\cap\partial D)\geq\delta$, where $\delta\leq\frac{1}{2}$ is the one in Lemma 4.1. Then by (ii) we get
$\delta^{p}\mu(B_{P,r}\cap D)$ $\leq\int_{B_{P,r}\cap D}\left(\omega_{x}^{D}(B_{P,\frac{r}{\delta}}\cap\partial D)\right)^{p}\,\mathrm{d}\mu(x)$
$=\int_{B_{P,r}\cap D}\left(\int_{\partial D}\chi_{B_{P,\frac{r}{\delta}}\cap\partial D}\,\mathrm{d}\omega_{x}^{D}\right)^{p}\,\mathrm{d}\mu(x)$
$\leq\int_{D}\left(\int_{\partial D}\chi_{B_{P,r}\cap\partial D}\,\mathrm{d}\omega_{x}^{D}\right)^{p}\,\mathrm{d}\mu(x)$
$\leq K\int_{\partial D}\left(\chi_{B_{P,r}\cap\partial D}\right)^{p}\,\mathrm{d}\omega^{D}_{O}$
$=K\omega_{O}^{D}(B_{P,r}\cap\partial D).$
Hence $\mu$ is a Carleson measure with its norm $M\leq\delta^{-p}K$. ∎
4.3. The proof of (i) $\Rightarrow$ (ii)
Let $D$ be a bounded Lipschitz domain in $\mathbb{R}^{n}$.
Proposition 4.2 (Harnack’s principle in Lipschitz domain (the inner cone case)).
Let $\Gamma$ be an inner cone with its vertex at $Q\in\partial D$ and $\widetilde{D}$ be a compact subset in $D\cap\Gamma$. Suppose the harmonic measure system satisfies (H). Then there is a constant $C=C(\widetilde{D},\Gamma)$ such that $H(\nu(\widetilde{D}),D)\leq C$ for any $k>0$ and $Q^{\prime}\in\partial D$ provided $\nu(D\cap\Gamma)\subset D$, where $\nu(x):=Q^{\prime}+k(x-Q)$.
Lemma 4.3.
If $r$ and $r_{0}$ with $r\geq r_{0}>0$ are sufficiently small, depending on the geometry of D, then for any $Q\in\partial D$ and any $P\in D\backslash C_{Q,r}$,
$\omega^{D}_{P}(\Delta_{Q,r_{0}})\leq C\omega^{D}_{T_{Q,r}}(\Delta_{Q,r_{0}}),$
where $C$ is a constant depending only on $D$.
Proof.
Denote $\Delta_{r}:=\Delta_{Q,r}=C_{Q,r}\cap\partial D$ and $T_{r}:=T_{Q,r}$. Then by Lemma 4.2 we get
$\omega^{D}_{T_{\delta r_{0}}}(\Delta_{r_{0}})\geq\delta,$
for $0<\delta\leq\frac{1}{2}$, where $\delta$ only depends on $D$. By the Harnack’s principle,
$\omega^{D}_{T_{r_{0}}}(\Delta_{r_{0}})\geq c_{0},$
where $c_{0}$ depends on $\delta$ and $D$, hence only depends on $D$. We can choose a constant $c_{1}=\frac{1}{c_{0}}$ such that
$\omega_{P}^{D}(\Delta_{r_{0}})\leq 1\leq c_{1}\omega_{T_{r_{0}}}^{D}(\Delta_{r_{0}})$
for any $P\in D\backslash C_{r_{0}}$. Suppose that for some positive integer $j$ with $2^{j}r_{0}\leq r$,
$\omega_{P}^{D}(\Delta_{r_{0}})\leq c_{1}\omega_{T_{2^{j-1}r_{0}}}^{D}(\Delta_{r_{0}})$
for any $P\in D\backslash C_{2^{j-1}r_{0}}$.
Let $P_{2}:=(X_{2},0)$, $\Omega=\{(X,y):|X|>1,y>-B|X-X_{2}|,\sqrt{X^{2}+y^{2}}<\frac{diam(D)}{r_{0}}\}$ where $X_{2}=(2,0,\cdots,0)$ in the standard coordinate system. Let $k(P)$ be the harmonic measure in $\Omega$ on the cylinder $\{(X,y)\in\partial\Omega:|X|=1\}$. For fixed point $Q^{\prime}$ on $\partial C_{2^{j}r_{0}}\cap\partial D$, scale $\Omega$ by the factor $2^{j-1}r_{0}$, then translate it such that the $y$-axis coincides with $OQ$ and $P_{2}$ becomes $Q^{\prime}$. We assume that the constant $B$ has been chosen depend on $D$ so that the cone part of $\partial\Omega$ is contained in $\mathbb{R}^{n}\backslash D$. Let $\Omega_{j}$ denote the region $\Omega$ scaled by the factor $2^{j-1}r_{0}$ and let $k_{j}$ be the corresponding harmonic measure $k$. By the maximum principle,
$\omega_{P}^{D}(\Delta_{r_{0}})\leq k_{j}(P)$
for any $P\in D\cap\Omega_{j}$. And by the strong maximum principle
$\omega_{P}^{D}(\Delta_{r_{0}})\leq c_{1}\omega_{T_{2^{j-1}r_{0}}}^{D}(\Delta_{r_{0}})k_{j}(P)$
for any $P\in\partial(D\cap\Omega_{j})$, hence for any $P\in D\cap\Omega_{j}$ by the maximum principle.
By the Harnack’s principle, there is a constant $c_{2}$ independent of $j$ such that
$\omega_{T_{2^{j-1}r_{0}}}^{D}(\Delta_{r_{0}})\leq c_{2}\omega_{T_{2^{j}r_{0}}}^{D}(\Delta_{r_{0}}).$
Since $k(P)\to 0$ as $\Omega\ni P\to P_{2}$, there is an $\eta>0$ such that $k(P)\leq\frac{1}{c_{2}}$ if $|P-P_{2}|<\eta$. Hence
(4.1) $\omega_{P}^{D}(\Delta_{r_{0}})\leq c_{1}\omega_{T_{2^{j}r_{0}}}^{D}(\Delta_{r_{0}})$
for any $P\in D\cap\Omega_{j}$ with $|P-Q^{\prime}|\leq\eta 2^{j-1}r_{0}$, hence for any $P\in D\cap\partial C_{2^{j}r_{0}}$ with $\text{dist}(P,\partial D\cap\partial C_{2^{j}r_{0}})\leq\eta 2^{j-1}r_{0}$. At the same time, by the Harnack’s inequality there is a constant $c_{3}$ such that
(4.2) $\omega_{P}^{D}(\Delta_{r_{0}})\leq c_{3}\omega_{T_{2^{j}r_{0}}}^{D}(\Delta_{r_{0}})$
for $P\in D\cap\partial C_{2^{j}r_{0}}$ with $\text{dist}(P,\partial D\cap\partial C_{2^{j}r_{0}})>\eta 2^{j-1}r_{0}$. Take $c_{4}=\max\{c_{1},c_{3}\}$. By (4.1) and (4.2) we get
$\omega_{P}^{D}(\Delta_{r_{0}})\leq c_{4}\omega_{T_{2^{j}r_{0}}}^{D}(\Delta_{r_{0}})$
for any $P\in D\cap\partial C_{2^{j}r_{0}}$, hence for any $P\in D\backslash C_{2^{j}r_{0}}$ by the maximum principle.
Since $r\geq r_{0}$ there exists a positive integer $j_{0}$ such that $2^{j_{0}}r_{0}\geq r\geq 2^{j_{0}-1}r_{0}$ the conclusion hods by the above inequality and Harnack principle. ∎
Lemma 4.4.
Let $E\subseteq\Delta:=B_{Q,r}\cap\partial D$ and $A:=A_{Q,r}$ be the point on the segment $OQ$ whose distance from $Q$ is $r$. Then
$\omega^{D}_{A}(E)\leq c\frac{\omega^{D}_{O}(E)}{\omega^{D}_{O}(\Delta)},$
where $c>0$ depends only on $D$.
Proof.
Consider a cone $\Gamma_{1}$ with vertex at the origin $O$, axis along $OQ$ and the disc $\Delta$ is contained in $\Gamma_{1}$. Let $H:=H(r,Q):=\Gamma_{1}\cap B_{Q,hr}$, where $h$ is a fixed constant. Let $\alpha$ denote the part of $\partial H$ which is contained in the standard inner cone $\Gamma_{2}$ at $Q$, and let $\beta$ denote the rest of $\partial H$ in $D$. For $h$ sufficiently large, depend on the Lipschitz constant of $D$, we may assume that $O\not\in B_{Q,hr}$ and there is a $t=t(h),0<t<1$, such that $tP\in\alpha$ for each $P\in\beta$. Let $\tilde{D}=D\backslash H$. By the maximum principle,
(4.3) $\omega_{P}^{D}(\Delta)\leq\omega_{P}^{\tilde{D}}(\alpha)+\omega_{P}^{\tilde{D}}(\beta)\leq(1+t)\omega_{P}^{\tilde{D}}(\alpha)$
for any $P\in\tilde{D}$.
Harnack’s principle implies
$\omega_{P}^{D}(E)\geq c\omega_{A}^{D}(E)$
for any $P\in\alpha$, where $c$ is independent of $r$. Hence, by the maximum principle,
(4.4) $\omega_{P}^{D}(E)\geq c\omega_{A}^{D}(E)\omega_{P}^{\tilde{D}}(\alpha)$
for any $P\in\tilde{D}$.
Setting $P=O$ in (4.3) and (4.4), we obtain
$\omega_{O}^{D}(E)\geq c\omega_{A}^{D}(E)\omega_{O}^{\tilde{D}}(\alpha)\geq\frac{c}{1+t}\omega_{A}^{D}(E)\omega_{O}^{D}(\Delta)\geq\frac{c}{2}\omega_{A}^{D}(E)\omega_{O}^{D}(\Delta),$
which proves the conclusion. ∎
For any domain $D$, a cone $\Gamma$ with vertex $Q\in\partial D$ is called a nontangential cone at $Q$ if there exists a cone $\widetilde{\Gamma}$ such that $\bar{\Gamma}\backslash\{Q\}\subset\widetilde{\Gamma}\subset D$. Denote $k(P,Q):=k^{D}_{O}(P,Q):=\frac{\,\mathrm{d}\omega_{P}^{D}(Q)}{\,\mathrm{d}\omega_{O}^{D}(Q)}$, for all $P\in D,Q\in\partial D$.
Lemma 4.5.
Let $\Gamma$ be any nontangential cone in $D$ at $Q_{0}$. For any $A^{\prime}\in\Gamma\cap D$ with $\mathrm{dist}(A^{\prime},Q_{0})=a$, denote $\Delta_{j}:=\Delta_{Q_{0},2^{j}a}$, $j=0,\cdots,N$, $R_{0}:=\Delta_{0}$ and $R_{j}:=\Delta_{j}\backslash\Delta_{j-1},\,j=1,\cdots,N$. Then
$\mathop{\mathrm{ess}\sup}\limits_{Q\in R_{j}}k(A^{\prime},Q)\leq\frac{cc_{j}}{\omega^{D}_{O}(\Delta_{j})},\,j=0,\cdots,N,$
with $\sum_{j=0}^{N}c_{j}\leq c^{\prime}<\infty$, where $c^{\prime}$ depends on $D$ and $c$ depends on $D$ and the shape of $\Gamma$.
Proof.
Let $A$ be the point on the segment $OQ_{0}$ whose distance from $Q_{0}$ is $a$. By Harnack’s priciple $\omega^{D}_{A^{\prime}}(E)\leq c\omega^{D}_{A}(E)$ where $c$ depends only on $D$ and the shape of $\Gamma$. Hence it is enough to prove the lemma if we show
(4.5) $\frac{\omega^{D}_{A}(E)}{\omega^{D}_{O}(E)}\leq\frac{cc_{j}}{\omega^{D}_{O}(\Delta_{j})}$
for arbitrarily small discs $E\subset R_{j}$.
Let $A_{j}$ be the point on the segment $OQ_{0}$ whose distance from $Q_{0}$ is $2^{j}a$, then by Lemma 4.4 we get
$\omega_{A_{j}}(E)\leq c\frac{\omega^{D}_{O}(E)}{\omega^{D}_{O}(\Delta_{j})}$
for any $E\subset\Delta_{j}$, where $c$ depends only on $D$. Suppose $E$ is small disc with center $Q_{j}\in R_{j},j\geq 4$. By Lemma 4.3, we get
$\omega^{D}_{P}(E)\leq c\omega^{D}_{T_{Q_{j},2^{j-2}a}}(E)$
for any $P\in D\backslash C_{Q_{j},2^{j-2}a}$. Harnack’s principle implies
$\omega^{D}_{T_{Q_{j},2^{j-2}a}}(E)\leq c\omega^{D}_{A_{j}}(E).$
Hence
$\omega^{D}_{P}(E)\leq c\frac{\omega^{D}_{O}(E)}{\omega^{D}_{O}(\Delta_{j})}$
for any $P\in D\backslash C_{Q_{j},2^{j-2}a}$.
Let
$\Sigma:=\{(X,y)\in\mathbb{R}^{n}:|X|^{2}+y^{2}<1,y>-B|X|\}$
and $h(P)$ be the harmonic measure of $\{(X,y)\in\partial\Sigma:|X|^{2}+y^{2}=1\}$. Denote $\Sigma_{j}$ be the region obtained by scaling $\Sigma$ by the factor $2^{j-2}a$ and translating it such that the $y$-axis coincides with $OQ_{0}$ and the origin coincides with $Q_{0}$. Denote $h_{j}$ be the harmonic measure corresponding to $h$. We assume $B$ has been chosen so that the standard outer cone at $Q_{0}$ coincide with the cone $\{y<-B|X|\}$. Since $D\cap\Sigma_{j}\subset D\backslash C_{Q_{j},2^{j-2}a}$, and the maximum principle imply
$\omega^{D}_{P}(E)\leq c\frac{\omega^{D}_{O}(E)}{\omega^{D}_{O}(\Delta_{j})}h_{j}(P)$
for $P\in D\cap\Sigma_{j}$. Taking $P=A$ we get
$\frac{\omega^{D}_{A}(E)}{\omega^{D}_{O}(E)}\leq\frac{ch_{j}(A)}{\omega^{D}_{O}(\Delta_{j})}.$
For $j=0,1,2,3$, we can choose $c_{j}$ depend on $D$ by the Harnack’s principle; For $j=4,\cdots,N$, we can choose $c_{j}:=h_{j}(A)$. Let $r=\max\{h(P):P\in\Sigma,|P|=\frac{1}{2}\}$, then $10<r<$ 1. by the maximum principle,
$h(\frac{P}{2})\leq rh(P)$
for any $P\in\Sigma$. Hence for $j\geq 4$,
$h_{j}(A)=h((0,\frac{1}{2^{j-2}}))\leq rh((0,\frac{1}{2^{j-3}}))=rh_{j-1}(A)\leq\cdots\leq r^{j-4}h_{4}(A).$
Since $10<r<$ 1 we get
$\sum_{j=4}^{N}c_{j}=\sum_{j=4}^{N}h_{j}(A)\leq\sum_{j=4}^{N}r^{j-4}h_{4}(A)\leq\frac{1}{1-r}<\infty.$
Then (4.5) is proved, which implies the conclusion of the lemma. ∎
Lemma 4.6.
Denote $\Delta:=\Delta_{Q_{0},r_{0}}$ for $Q_{0}\in\partial D$. Then
$\mathop{\mathrm{ess}\sup}\limits_{Q^{\prime}\in\partial D\backslash\Delta}k(P,Q^{\prime})\to 0,$
as $P\to Q_{0}$.
Proof.
Denote $\Sigma_{0}$ be the region by Scaling $\Sigma$ in the proof of lemma 4.5 by the factor $\frac{r_{0}}{3}$ and translating it to $Q_{0}$ as in the proof of lemma 4.5. Denote $h_{0}$ be the harmonic measure corresponding to $h$ defined in the proof of Lemma 4.5. For any small disc $\Delta^{\prime}:=\Delta^{\prime}_{Q^{\prime},r^{\prime}}$ with $Q^{\prime}\in\partial D\backslash\Delta$ and $0<r^{\prime}<\frac{r_{0}}{3}$, Lemma 4.3 implies
$\omega^{D}_{P}(\Delta^{\prime})\leq c_{1}\omega^{D}_{T_{Q^{\prime},\frac{r_{0}}{3}}}(\Delta^{\prime})$
for any $P\in D\backslash C_{Q^{\prime},\frac{r_{0}}{3}}$, where $c_{1}$ depends only on $D$. By Harnack inequality, we get
$\omega^{D}_{T_{Q^{\prime},\frac{r_{0}}{3}}}(\Delta^{\prime})\leq c_{2}\omega^{D}_{O}(\Delta^{\prime}),$
where $c_{2}$ depends on $D$ and $r_{0}$. Then the above two inequalities imply
$\omega^{D}_{P}(\Delta^{\prime})\leq c_{1}c_{2}\omega^{D}_{O}(\Delta^{\prime})$
for any $P\in\partial\Sigma_{0}\cap D\subseteq D\backslash C_{Q^{\prime},\frac{r_{0}}{3}}$. Since $0\leq\omega_{P}(\Delta^{\prime})\leq 1=h_{0}(P)$ for $P\in\partial\Sigma_{0}\cap D$ and $\omega_{P}(\Delta^{\prime})=0\leq h_{0}(P)$ for $P\in\partial D\cap\Sigma_{0}$, the maximum principle implies
$\omega^{D}_{P}(\Delta^{\prime})\leq h_{0}(P),$
for any $P\in\Sigma_{0}\cap D$. Hence
$\omega^{D}_{P}(\Delta^{\prime})\leq c_{1}c_{2}\omega^{D}_{O}(\Delta^{\prime})h_{0}(P)$
for any $P\in\partial(\Sigma_{0}\cap D)$, then for $P\in\Sigma_{0}\cap D$ by the maximum principle. Since $h_{0}(P)\to 0$ as $P$ tends to $Q_{0}$, we have
$\frac{\omega^{D}_{P}(\Delta^{\prime})}{\omega^{D}_{O}(\Delta^{\prime})}\to 0$
as $P\to Q_{0}$ uniformly for any radius small $\Delta^{\prime}\subset\partial D\backslash D$ containing $Q^{\prime}$. This proves the conclusion by the definition of $k(P,Q^{\prime})$. ∎
We need to prove the following lemma.
Lemma 4.7.
Suppose $D\subset\mathbb{R}^{n}(n\geq 3)$ is a Lipschitz domain, starlike about the origin $O\in D$ and $\mu$ is a Carlson measure in $D$. For a $Q_{0}\in\partial D$ and appropriate $A,B>0$ depend on $D$, denote $S_{1}:=\{Q\in\partial D:\mathrm{dist}(Q,l_{OQ_{0}})<B\}$, $D_{2}:=\{P\in D:\mathrm{dist}(P,Q_{0})<2A,\mathrm{dist}(P,l_{OQ_{0}})<2B)\}$. If there exists a $\delta>0$ such that for any $x\in D$, $\omega_{x}^{D}\in B_{1+\delta}(\omega_{O}^{D})$, then for any $f\in L^{p}(\partial D,\omega_{\cdot}^{D})$, there exists $K>0$ such that
[…]
where $p=(1+\delta)^{\prime}$.
Proof.
It is sufficient to prove the lemma in the case $f\geq 0$, since $\mu\{P\in D_{2}:|\omega^{D}_{P}(f)|>s\}\leq\mu\{P\in D_{2}:|\omega^{D}_{P}(f^{+})|>\frac{s}{2}\}+\mu\{P\in D_{2}:|\omega^{D}_{P}(f^{-})|>\frac{s}{2}\}$.
Denote $S_{i}:=\{Q\in\partial D:\mathrm{dist}(Q,l_{OQ_{0}})<iB\}$ and $D_{i}:=\{P\in D:\mathrm{dist}(P,Q_{0})<iA,\mathrm{dist}(P,l_{OQ_{0}})<iB\}$ for $i\in\mathbb{Z}^{+}$. For any $Q\in S_{3}$ we can choose $\Gamma(Q)$ being a small nontangential inner cone with center line paralleling to $OQ_{0}$ and the height $h>0$ small enough such that:
  1. (1)
    For any $Q\in S_{2}$, $\Gamma(Q)\subseteq D_{3}$ and
  2. (2)
    by Lemma 4.6, we can choose $\Delta_{h}:=\Delta_{Q,r_{h}}$ such that for any $Q\in S_{3}$ and any $P^{\prime}\in\Gamma(Q)$,
    $\mathop{\mathrm{ess}\sup}\limits_{Q^{\prime}\in\partial D\backslash\Delta_{h}}k(P^{\prime},Q^{\prime})\leq cc^{\prime},$
    where $c,c^{\prime}$ are the constants in Lemma 4.5.
Then Lemma 4.5 implies that for any $Q\in S_{3}$ and any $P^{\prime}\in\Gamma(Q)$,
$\omega^{D}_{P^{\prime}}(f)$ $=\int_{\Delta_{h}}f\,\mathrm{d}\omega^{D}_{P^{\prime}}+\int_{\partial D\backslash\Delta_{h}}f\,\mathrm{d}\omega^{D}_{P^{\prime}}$
$\leq\sum_{j=0}^{m(|P^{\prime}Q|)}\int_{R_{j}}fk(P^{\prime},\cdot)\,\mathrm{d}\omega^{D}_{O}+\int_{\partial D\backslash\Delta_{h}}fk(P^{\prime},\cdot)\,\mathrm{d}\omega^{D}_{O}$
$\leq c{\sum_{j=0}^{\infty}{c_{j}}\fint_{\Delta_{j}}f\,\mathrm{d}\omega^{D}_{O}}+\mathop{\mathrm{ess}\sup}\limits_{Q^{\prime}\in\partial D\backslash\Delta_{h}}k(P^{\prime},Q^{\prime})\fint_{\partial D}f\,\mathrm{d}\omega^{D}_{O}$
$\leq c\sum_{j=0}^{\infty}{c_{j}}f^{*}(Q)+cc^{\prime}f^{*}(Q)$
$=2cc^{\prime}f^{*}(Q),$
where $c^{\prime}$ depends only on $D$, $c$ depends on $D$ and the shape of $\Gamma(Q)$, hence depends only on $D$ and
$f^{*}(Q):=\sup\limits_{r>0}\fint_{\Delta(Q,r)}f\,\mathrm{d}\omega^{D}_{O}.$
Then we get for any $Q\in S_{3}$,
(4.6) $\sup_{P^{\prime}\in\Gamma(Q)}\omega^{D}_{P^{\prime}}(f)\leq C_{0}f^{*}(Q),$
where $C_{0}$ depends only on $D$.
We divide the rest of proof into two steps.
Step 1. Let $V:=D_{2}\cap(\mathop{\cup}\limits_{Q\in S_{2}}\Gamma(Q))$, $U:=D_{2}-V$. By the Harnack inequality and Hölder inequality, $\omega^{D}_{P}(f)\leq C\omega^{D}_{O}(f)\leq C\omega^{D}_{O}(f^{p})$ for any $P\in U$ and $f\in L^{p}(\partial D,\omega_{\cdot}^{D})$. Hence, for $s\geq C^{\frac{1}{p}}\|f\|_{p,\omega_{O}^{D}}$, we get $C\omega^{D}_{O}(\left(\frac{f}{s}\right)^{p})\leq 1$, which implies $\omega^{D}_{P}(\frac{f}{s})\leq 1$, i.e.
$\mu\{P\in U:\omega^{D}_{P}(f)>s\}=\mu\{P\in U:\omega^{D}_{P}(\frac{f}{s})>1\}=0\leq Cs^{-p}\|f\|_{p,\omega^{D}_{O}}^{p};$
And for $0<s<C^{\frac{1}{p}}\|f\|_{p,\omega_{O}^{D}}$, which implies $1<Cs^{-p}\|f\|_{p,\omega^{D}_{O}}^{p}$, then
$\mu\{P\in U:\omega^{D}_{P}(f)>s\}\leq\mu(D)\cdot 1<\mu(D)Cs^{-p}\|f\|_{p,\omega^{D}_{O}}^{p}.$
Then the above two inequalities imply for any $s>0$,
(4.7) $\mu\{P\in U:\omega^{D}_{P}(f)>s\}\leq C^{\prime}s^{-p}\|f\|_{p,\omega^{D}_{O}}^{p},$
where $C^{\prime}:=\max\{C,\mu(D)C\}$ depends only on $D$.
Step 2. We will prove
(4.8) $\mu\{P\in V:\omega_{P}^{D}(f)>s\}\leq Cs^{-p}\|f\|_{p,\omega_{O}^{D}}^{p}.$
[…]
Suppose $F\subset\{P\in V:\omega_{P}^{D}(f)>s\}$ is compact. Then $F\subset\cup_{P\in F}R(P)$. Denote the width of $R(P)$ by $r(P):=t\mathrm{dist}(P,P^{*})$ and $r_{0}:=\mathop{\sup}\limits_{P\in F}r(P)$. Then $r_{0}<th$. Select $P_{1}\in F$ such that $r(P_{1})>\frac{r_{0}}{2}$. Define
$F_{1}:=F\backslash R(P_{1}),\,r_{1}:=\sup_{P\in F_{1}}\{r(P)\},$
and select $P_{2}\in F_{1}$ such that $r(P_{2})>\frac{r_{1}}{2}$. Continue this process until the first integer $m$ is found such that $F_{m}$ is an empty set. If no such integer exists, continue this process and set $m=+\infty$ and $F\backslash\cup_{i=1}^{\infty}R(P_{i})\not=\emptyset$.
Suppose that $i>j$. Then $P_{i}\in F_{i-1}\subset F\backslash R(P_{j})$, thus $P_{i}\notin R(P_{j})$ and $r(P_{i})\leq r_{i-1}<2r(P_{j})$. It easily follows that $\frac{1}{3}R(P_{i})\cap\frac{1}{3}R(P_{j})=\emptyset$ by $P_{i}\notin R(P_{j})$, where $\frac{1}{3}R(P):=R(\frac{2}{3}P^{*}+\frac{1}{3}P)$. If $m=+\infty$, then there exists a subsequence of $\{r(P_{i})\}$ converging to zero since $\frac{1}{3}R(P_{i})$ are pairwise disjoint and have centers in a bounded set $\frac{1}{3}F\subset V$. In this case since $F\backslash\cup_{i=1}^{\infty}R(P_{i})\not=\emptyset$, for any $P\in F\backslash\cup_{i=1}^{\infty}R(P_{i})$, then $P\in F_{i}$ and $r(P)>r(P_{i})$ for all $i\in\mathbb{N}$, a contradiction. Thus $m<\infty$.
Fix a $y\in\mathbb{R}^{n}$, then $\mathbb{R}^{n}$ can be decomposed into $2^{n}$ parts by $n$ hyperplanes which are parallel to coordinate hyperplanes and pass through the point $y$. Choose one of the $2^{n}$ part and denote it by $O$. Set $|z|_{\infty}:=\max\{\sup_{1\leq i\leq n-1}|z^{i}|,t|z^{n}|\}$ for $z=(z^{1},\cdots,z^{n})\in\mathbb{R}^{n}$. We pick a $P_{k_{0}}\in F\cap O$ such that $y\in R(P_{k_{0}})$ and $|P_{k_{0}}-y|_{\infty}$ is the largest one among all $P_{j}\in F\cap O$. Then we claim that $P_{j}\in R(P_{k_{0}})$ for any $P_{j}\in F\cap O$ with $y\in R(P_{j})$. To show this we notice that
$|P_{j}^{i}-P_{k_{0}}^{i}|$ $=||P_{j}^{i}-y^{i}|-|P_{k_{0}}^{i}-y^{i}||$
$\leq\max(|P_{k_{0}}-y|_{\infty},|P_{j}-y|_{\infty})$
$=|P_{k_{0}}-y|_{\infty}<\frac{1}{2}r(P_{k_{0}}),$
holds for each $i=1,\cdots,n$. And we have $j\leq k_{0}$ when $P_{j}\in R(P_{k_{0}})$ by previous proof, which implies that $\alpha=\frac{1}{2}r(P_{k_{0}})<r(P_{j})$. Let $\{R(P_{k_{i}})\}_{i\in I}$ be those cylinders with $\frac{1}{2}\alpha<r(P_{k_{0}})<\alpha$. Then we have
$\frac{\alpha^{n}|I|}{3^{n}}\leq\sum_{i\in I}|\frac{1}{3}R(P_{k_{i}})|\leq|\cup_{i\in I}R(P_{k_{i}})|\leq(4\alpha)^{n}.$
This shows that $|I|\leq 12^{n}$, and since there are $2^{n}$ sets $O$, there are finitely many $P_{j}\in F$ such that $F\subset\cup R(P_{j})$ and $\sum\chi_{R(P_{j})}\leq C_{1}$, where $C_{1}=24^{n}$ depends only on $n$.
Hence $\mu(F)\leq\sum\mu(R(P_{j}))\leq\sum M\omega^{D}_{O}(R(P_{j})\cap\partial D)\leq C_{1}M\omega^{D}_{O}\{Q\in S_{3}:f^{*}(Q)>\frac{1}{C_{0}}s\}$. Since $F$ is arbitrary, it follows that
(4.9) $\mu\{P\in V:\omega^{D}_{P}(f)>s\}\leq C\omega^{D}_{O}\{Q\in S_{3}:f^{*}(Q)>\frac{1}{C_{0}}s\}$
The lemma follows if we can show that $\omega_{O}^{D}\{Q\in S_{3}:f^{*}(Q)>\frac{1}{C_{0}}s\}\leq Cs^{-p}\|f\|^{p}_{p,\omega_{O}^{D}}$. In fact, by the $B_{1+\delta}$ condition, for $Q\in S_{3}$, we get
[…]
Hence
[…]
and
[…]
This implies
(4.10) $\omega_{O}^{D}\{Q\in S_{3}:f^{*}(Q)>\frac{1}{C_{0}}s\}\leq Cs^{-p}\|f\|_{p,\omega_{O}^{D}}^{p}.$
The proof of (i) $\Rightarrow$ (ii).
The conclusion holds by the Marcinkiewicz interpolation theorem and the boundedness of $D$. ∎
5. Acknowledgments
This research is partly supported by the National Natural Science Foundation of China (NSFC 11501268 and 11471147) and the Fundamental Research Funds for the Central Universities (lzujbky-2016-101).